# Quantile Reward Policy Optimization: Alignment with Pointwise Regression and Exact Partition Functions

### v1

@@ Appendix C The partition function is a moment generating function at $t=1/{\beta}$
$=\sum_{r}\exp\Bigl{(}\tfrac{1}{\beta}r\Bigr{)}\,\underbrace{\Pr_{y\sim\pi_{\mathrm{ref}}(\cdot\mid x)}\{\mathcal{R}(x,y)=r\}}_{P_{\mathrm{ref}}(r\mid x)}$ [The inner sum becomes a cumulative probability of all $y$’s with the reward $r$Rightarrow it’s a probability of obtaining the reward $r$]
@@
Appendix D Distribution of the quantile reward $\mathcal{R}_{q}$
@@
Appendix E Partition function for the quantile reward $\mathcal{R}_{q}$ and the generalized transformed reward $\tilde{\mathcal{R}}$
Here we compute the partition function $\tilde{Z}$ for the case of a generalized transformed reward ~$\tilde{\mathcal{R}}(x,y)=f(\mathcal{R}_{q}(x,y))$ (as introduced in Equation 13). To get $Z_{q}$ for the quantile reward $\mathcal{R}_{q}$, one can just take $f(t)=t$.
@@
~$\tilde{\mathcal{R}}(x,y)=f(\mathbf{\mathcal{R}}_{q}(x,y))$
@@
Following the results obtained in Appendix C, we can express the partition function in terms of the moment generating function (MGF) for the generalized transformed reward ~$\tilde{\mathcal{R}}(x,y)$ distribution for $y\sim\pi_{ref}(\cdot\mid x)$:
@@
where, similar to Appendix C, we introduce $\tilde{P}_{\mathrm{ref}}(\cdot\mid x)$—the probability distribution of the generalized transformed reward ~$\tilde{\mathcal{R}}(x,y)$ induced by $y\sim\pi_{ref}(\cdot\mid x)$.
If one knows the distribution of the transformed reward, the partition function can be computed directly by evaluating its MGF at $t=1/\beta$. For instance, in Appendix D, we show that the quantile reward $\mathcal{R}_{q}(x,y)$, when treated as a continuous variable, follows a $\mathrm{Uniform}(0,1)$ distribution. The MGF for $\mathrm{Uniform}(0,1)$ distribution is given by $M_{r_{q}}(t)=\frac{1}{t}(e^{t}-1)$, and thus the corresponding partition function is
@@
[…]
@@
$\tilde{Z}(x)=\mathbb{E}_{\tilde{r}\sim\tilde{P}_{\mathrm{ref}}(\cdot\mid x)}\left[\exp\Bigl{(}\tfrac{1}{\beta}\tilde{r}\Bigr{)}\right]$ $=\mathbb{E}_{r\sim P_{ref}(\cdot\mid x)}\left[\exp\Bigl{(}\tfrac{1}{\beta}f(F_{ref}(x,r)))\Bigr{)}\right]$
@@
Here, $P_{\mathrm{ref}}(\cdot\mid x)$ denotes the probability distribution of the initial reward $\mathcal{R}(x,y)$ induced by $y\sim\pi_{ref}(\cdot\mid x)$, as introduced in Appendix C and similar to $\tilde{P}_{\mathrm{ref}}(\cdot\mid x)$ introduced earlier in this section; $\hat{P}_{\mathrm{ref}}(\cdot\mid x)$ is an empirical estimate for $P_{\mathrm{ref}}(\cdot\mid x)$; $F_{ref}(x,r)$ is a CDF of initial rewards under the reference policy as introduced in Definition 8. $K$ denotes the number of possible initial reward values.
The expression for the discrete case derived above is intended to be used in the case when the number of possible reward values is small (i.e. $<10$). For other cases, the continuous version is likely to be a decent approximation for the partition function, and is more preferable, since it does not require one to estimate the probabilities ${\hat{P}_{\mathrm{ref}}(r_{i}\mid x)}_{i}$.
Appendix F Sufficient conditions for the function in the reward transformation for the existence of a finite partition function
@@
Recall that the partition function can be expressed as a moment generating function for the distribution of the reward used in the objective 1 (see Appendix C for the derivation). Here we consider objective 1 with the generalized transformed reward $\tilde{\mathcal{R}}$:
$\tilde{Z}=\mathbb{E}_{\tilde{r}\sim\tilde{P}_{\mathrm{ref}}(\cdot\mid x)}\left[\exp\Bigl{(}\tfrac{1}{\beta}\tilde{r}\Bigr{)}\right],$
where $\tilde{P}_{\mathrm{ref}}(r\mid x)$ denotes the probability distribution of the generalized transformed reward ~$\tilde{\mathcal{R}}(x,y)$ induced by the reference policy $y\sim\pi_{ref}(\cdot\mid x)$.
@@
Appendix G Noise tradeoff in the calibrated target: estimating the partition function vs. estimating the quantile reward
@@
Direct regression methods described above require estimating a partition function $Z$ that is computationally very expensive. To support this empirical observation, we provide a theoretical analysis of the noise amplitude in its empirical estimate, assuming that completion rewards for a given prompt follow a Gaussian distribution. The standard deviation of a logarithm of an empirical estimate of partition function $\hat{Z}$ has the following expression for a large number of samples $n\gg e^{\frac{\sigma^{2}}{\beta^{2}}}$ (derivation in Appendix H):
@@
To assess the impact of this change, we compare the noise level in the objective 6 after applying the transformation with the initial noise level. The standard deviation of $\beta\log{\hat{Z}(x)}$ for the initial reward was previously shown in Equation 17. The standard deviation of $\mathbf{\mathcal{R}}_{q}(x,y)$ is given by the following expression (proof in Appendix I):
@@
Appendix H Noise amplitude in the partition function estimate for the initial reward
Here we compute the variance of the $\log{\hat{Z}(x)}$ estimate for the general case of using the initial reward for the policy fitting regression objective. Recall that the partition function can be expressed as a moment generating function for the distribution of the reward used in the objective 1 (see Appendix C for the derivation):
$Z=\mathbb{E}_{r\sim P_{\mathrm{ref}}(\cdot\mid x)}\left[\exp\Bigl{(}\tfrac{1}{\beta}r\Bigr{)}\right],$
@@ Appendix H Noise amplitude in the partition function estimate for the initial reward
Appendix I Noise amplitude in the transformed reward estimate
@@
Appendix J Equivalence of all quantile transformation functions strictly increasing and upper-bounded on the interval $[0,1]$
[…]
@@
~$\tilde{\mathcal{R}}(x,y)=f\left(\mathbf{\mathcal{R}}_{q}(x,y)\right),$
@@
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\left\{\frac{1}{\beta}f\left(F_{ref}(x,r)\right)-\frac{1}{\beta}g\left(F_{ref}(x,r)\right)+\log{\frac{Z_{g}(x)}{Z_{f}(x)}}\right\}$
@@
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\underbrace{\sum_{y:\mathcal{R}(x,y)=r}\pi_{ref}(y\mid x)}_{P_{ref}(r|x)}$
@@
where we used the fact that $f(1)=g(1)=0$ and $f^{\prime}_{-}(1)=g^{\prime}_{-}(1)=1$ after rescale and shift described in Appendix J.1.
@@
The partition function has the following asymptotics (see Appendix K for more details):
@@
Appendix K Asymptotics of the partition function $Z$ for the family of quantile transformation functions strictly increasing and upper-bounded on the interval $[0,1]$
@@
We assume mild regularity conditions, specifically $f\in C^{\infty}[0,1]$. We also assume that $f$ is normalized according to the rescale and shift procedure described in Appendix J.1.
Recall that the partition function can be expressed as a moment generating function for the distribution of the reward used in the objective 1 (see Appendix C for the derivation). Here we consider objective 1 with the generalized transformed reward $\tilde{\mathcal{R}}$:
$\tilde{Z}=\mathbb{E}_{\tilde{r}\sim\tilde{P}_{\mathrm{ref}}(\cdot\mid x)}\left[\exp\Bigl{(}\tfrac{1}{\beta}\tilde{r}\Bigr{)}\right],$
where $\tilde{P}_{\mathrm{ref}}(r\mid x)$ denotes the probability distribution of the generalized transformed reward ~$\tilde{\mathcal{R}}(x,y)$ induced by the reference policy $y\sim\pi_{ref}(\cdot\mid x)$.
@@
$\tilde{Z}=\mathbb{E}_{\tilde{r}\sim\tilde{P}_{\mathrm{ref}}(\cdot\mid x)}$ $\left[\exp\Bigl{(}\tfrac{1}{\beta}\tilde{r}\Bigr{)}\right]=\mathbb{E}_{r_{q}\sim\mathrm{U}[0,1]}\left[\exp\Bigl{(}\tfrac{1}{\beta}f\bigl{(}r_{q}\bigl{)}\Bigr{)}\right]=\int_{0}^{1}\exp{\left(\frac{1}{\beta}f\left(z\right)\right)}~{}dz.$
Here we used the fact that the probability distribution of the quantile-reward $\mathcal{R}_{q}(x,y)$ induced by the reference policy $y\sim\pi_{ref}(\cdot\mid x)$ is uniform from 0 to 1 (see Appendix D for the derivation).
@@
Note, that we assume the function $f$ to be normalized according to the rescale and shift procedure described in Appendix J.1 (as we require this property in the transformation functions equivalence proof J). Without this normalization the asymptotics will have a different form which can be derived similarly using the Laplace’s method.
Appendix L Detailed experimental protocol
@@
Llama 8B Tülu 3 SFT (Lambert et al., 2024) Llama-3.1-8B Tülu 3 SFT dataset
Mistral 7B Instruct v0.2 (Jiang, 2024) Mistral-7B-v0.2 Unknown public instruction datasets
@@
Magpie-Air (Xu et al., 2024b) General Chat 98,000 2,000 Length $>$ 2048 tokens $<$ 1%
UltraFeedback (Cui et al., 2024) General Chat 61,135 2,000 Length $>$ 2048 tokens $<$ 1%
LeetCode (Xia et al., 2025) Coding 2,641 228 Length $>$ 2048 tokens or $<$ 20 test cases $<$ 3%
@@
General Chat ArmoRM (Wang et al., 2024) Reward model with sequences truncated at 2048 tokens
General Chat Length-controlled ArmoRM (new) Length bias canceled via regression as in AlpacaEval 2 (Dubois et al., 2024)
@@
We release a reference implementation of QRPO based on a HuggingFace TRL  (von Werra et al., 2020) trainer also supporting the baselines we trained (DPO, REBEL, SimPO) at https://github.com/CLAIRE-Labo/quantile-reward-policy-optimization. The codebase is based on the Python Machine Learning Research Template (Moalla, 2025). It contains
@@
L.2 Initial SFT-only Models
We consider popular middle-scale models that have been instruction fine-tuned (IFT/SFT) but have not undergone RL fine-tuning.555 This is important as our experiments emulate the RL fine-tuning stage after the supervised fine-tuning stage in a post-training pipeline, and more importantly because reporting improvements on only specific benchmarks over already (usually closed-source) RL fine-tuned models is a biased view on the performance profile of these models which have been carefully tuned to balance multiple benchmarks. E.g., taking Llama-3-8B-Instruct and improving its AlpacaEval 2 scores by RL fine-tuning it on a narrow chat dataset is not a faithful way to report an algorithmic improvement as it does not show where the model has regressed. We choose allenai/Llama-3.1-Tulu-3-8B-SFT (Llama 8B Tülu 3 SFT) (Lambert et al., 2024), a fine-tuning of the meta-llama/Llama-3.1-8B base model on the Tülu 3 SFT dataset (Lambert et al., 2024), and mistralai/Mistral-7B-Instruct-v0.2 (Mistral 7B Instruct v0.2) (Jiang, 2024).
L.3 Datasets
For the general chat task, we use two popular alignment datasets, extensively used in previous work (Meng et al., 2024; Gao et al., 2024) consisting primarily of synthetic data.
Magpie-Align/Magpie-Air-DPO-100K-v0.1 (Magpie-Air) (Xu et al., 2024b) is a strong alignment dataset containing 98000 training samples (and 2000 testing samples) with synthetic instructions and completions generated by Llama-3-8B-Instruct targeting mainly information seeking, and other categories such as creative writing, advice seeking, planning, and math. The chosen (resp. rejected) completions have been selected by picking the best (resp. worse) completions in a set of 5 completions rated by the RLHFlow/ArmoRM-Llama3-8B-v0.1 reward.
HuggingFaceH4/ultrafeedback_binarized (UltraFeedback) (Cui et al., 2024) consists of 61135 training samples (and 2000 testing samples) with instructions targeting instruction following, truthfulness, honesty, helpfulness, and completions generated by a large pool of models with different capabilities (between 7B and 70B open models and closed models including GPT4), annotated by GPT-4. This is the pre-processed version of the original UltraFeedback dataset that was used to train Zephyr-7B-$\beta$ (Tunstall et al., 2024), where for each prompt, the completion with the highest overall score is taken as the chosen completion, and one of the remaining three at random as the rejected one.
For the coding task, we use newfacade/LeetCodeDataset (LeetCode) (Xia et al., 2025), which contains 2641 training samples (and 228 testing samples) comprising easy, medium, and hard LeetCode problems. Each problem has a median of 100 test cases and contains a correct solution that we use for additional SFT before performing the RL fine-tuning phase.
@@
L.4 Rewards
For the general chat task we use the RLHFlow/ArmoRM-Llama3-8B-v0.1 (ArmoRM) (Wang et al., 2024) reward model with a maximum sequence length of $2048$. It is used to compute the rewards directly used by REBEL and QRPO in all distribution settings and to label the chosen and rejected preferences for DPO and SimPO in the off-policy setting. This also holds for Magpie-Air in the offline setting as its chosen and rejected annotations have been annotated with ArmoRM as well by Xu et al. (2024b). For UltraFeedback the offline chosen and rejected annotations are not annotated by the ArmoRM, but most of the best models for the baselines turned out to be from the off-policy setting due to the low quality of the UltraFeedback offline completions.
Although ArmoRM has a verbosity component to reduce length bias, it still suffers from out-of-distribution length bias. i.e., its length bias has been minimized on specific outputs (original UltraFeedback), but is still present when computing the rewards on chats that diverge from the UltraFeedback outputs, which we did observe in our experiments. We therefore also report a length-controlled reward whose coefficients are computed for each model based on a regression that cancels the length component as done in AlpacaEval 2 (Dubois et al., 2024) (see Appendix L.5 for more details on the length-controlled reward computation).
For LeetCode, we use a Python sandbox for executing LLM-generated code in a safe environment with exclusive and controlled hardware limits, which we built for this project. The reward for a solution to a problem is the average number of test cases passed in the suite of tests for the problem. Each problem has a minimum of 20 tests and a median of 100 tests, and we cap the number of tests to run for each problem to the first 100 tests for computational efficiency. We use this average pass rate reward instead of a binary reward indicating whether all tests passed for two main reasons: (i) because the return value of the solution for edge cases is often not specified, still those appear in the test cases with arbitrary targets causing high rewards to collapse when binarized (e.g., passing the test cases with a rate of 0.99 because of a single failed test on an under-specified edge case would be converted to a 0 binary reward). (ii) because we assume a reward that can be treated as continuous in the theory of QRPO.666 We have also derived a version of QRPO for rewards with a small number of possible values in Appendix E which would be compatible with the binary reward in this case. We defer this setting to future work.
L.5 Length-controlled reward (LC-Reward) computation

### v2

@@ Appendix D The partition function is a moment generating function at $t=1/{\beta}$
$=\sum_{r}\exp\Bigl(\tfrac{1}{\beta}r\Bigr)\,\underbrace{\Pr_{y\sim\pi_{\mathrm{ref}}(\cdot\mid x)}\{\mathcal{R}(x,y)=r\}}_{P_{\mathrm{ref}}(r\mid x)}$ [The inner sum becomes a cumulative probability of all $y$’s with the reward $r$ $\Rightarrow$ it’s a probability of obtaining the reward $r$]
@@
Appendix E Distribution of the quantile reward $\mathcal{R}_{q}$
@@
As an illustration, we show in Figure 8 an empirical distribution of the quantile reward.
Figure 8: Empirical distribution of the quantile reward $\mathcal{R}_{q}(x,y)$ for completions generated from a reference model $\pi_{ref}$.
Appendix F Partition function for the quantile reward $\mathcal{R}_{q}$ and the generalized transformed reward $\tilde{\mathcal{R}}$
Here we compute the partition function $\tilde{Z}$ for the case of a generalized transformed reward $\tilde{\mathcal{R}}(x,y)=f(\mathcal{R}_{q}(x,y))$ (as introduced in Equation 13). To get $Z_{q}$ for the quantile reward $\mathcal{R}_{q}$, one can just take $f(t)=t$.
@@
$\tilde{\mathcal{R}}(x,y)=f(\mathbf{\mathcal{R}}_{q}(x,y))$
@@
Following the results obtained in Appendix D, we can express the partition function in terms of the moment generating function (MGF) for the generalized transformed reward $\tilde{\mathcal{R}}(x,y)$ distribution for $y\sim\pi_{ref}(\cdot\mid x)$:
@@
where, similar to Appendix D, we introduce $\tilde{P}_{\mathrm{ref}}(\cdot\mid x)$—the probability distribution of the generalized transformed reward $\tilde{\mathcal{R}}(x,y)$ induced by $y\sim\pi_{ref}(\cdot\mid x)$.
If one knows the distribution of the transformed reward, the partition function can be computed directly by evaluating its MGF at $t=1/\beta$. For instance, in Appendix E, we show that the quantile reward $\mathcal{R}_{q}(x,y)$, when treated as a continuous variable, follows a $\mathrm{Uniform}(0,1)$ distribution. The MGF for $\mathrm{Uniform}(0,1)$ distribution is given by $M_{r_{q}}(t)=\frac{1}{t}(e^{t}-1)$, and thus the corresponding partition function is
@@
$\tilde{Z}=\mathbb{E}_{\tilde{r}\sim\tilde{P}_{\mathrm{ref}}(\cdot\mid x)}\left[\exp\Bigl(\tfrac{1}{\beta}\tilde{r}\Bigr)\right]=\mathbb{E}_{r_{q}\sim\mathrm{U}[0,1]}\left[\exp\Bigl(\tfrac{1}{\beta}f(r_{q})\Bigr)\right]=\int_{0}^{1}\exp\Bigl(\tfrac{1}{\beta}f(t)\Bigr)\,dt.$
@@
[…]
@@
Here, $P_{\mathrm{ref}}(\cdot\mid x)$ denotes the probability distribution of the initial reward $\mathcal{R}(x,y)$ induced by $y\sim\pi_{ref}(\cdot\mid x)$, as introduced in Appendix D and similar to $\tilde{P}_{\mathrm{ref}}(\cdot\mid x)$ introduced earlier in this section; $\hat{P}_{\mathrm{ref}}(\cdot\mid x)$ is an empirical estimate for $P_{\mathrm{ref}}(\cdot\mid x)$; $F_{ref}(x,r)$ is a CDF of initial rewards under the reference policy as introduced in Definition 8. $K$ denotes the number of possible initial reward values.
The expression for the discrete case derived above is intended to be used in the case when the number of possible reward values is small (i.e. <10<10). For other cases, the continuous version is likely to be a decent approximation for the partition function, and is more preferable, since it does not require one to estimate the probabilities ${\hat{P}_{\mathrm{ref}}(r_{i}\mid x)}_{i}$.
Appendix G Sufficient conditions for the function in the reward transformation for the existence of a finite partition function
@@
Recall that the partition function can be expressed as a moment generating function for the distribution of the reward used in the objective 1 (see Appendix D for the derivation). Here we consider objective 1 with the generalized transformed reward $\tilde{\mathcal{R}}$:
[…]
where $\tilde{P}_{\mathrm{ref}}(r\mid x)$ denotes the probability distribution of the generalized transformed reward $\tilde{\mathcal{R}}(x,y)$ induced by the reference policy $y\sim\pi_{ref}(\cdot\mid x)$.
@@
Appendix H Noise tradeoff in the calibrated target: estimating the partition function vs. estimating the quantile reward
@@
Direct regression methods described above require estimating a partition function $Z$ that is computationally very expensive. To support this empirical observation, we provide a theoretical analysis of the noise amplitude in its empirical estimate, assuming that completion rewards for a given prompt follow a Gaussian distribution. The standard deviation of a logarithm of an empirical estimate of partition function $\hat{Z}$ has the following expression for a large number of samples $n\gg e^{\frac{\sigma^{2}}{\beta^{2}}}$ (derivation in Appendix I):
@@
To assess the impact of this change, we compare the noise level in the objective 6 after applying the transformation with the initial noise level. The standard deviation of $\beta\log{\hat{Z}(x)}$ for the initial reward was previously shown in Equation 18. The standard deviation of $\mathbf{\mathcal{R}}_{q}(x,y)$ is given by the following expression (proof in Appendix J):
@@
Appendix I Noise amplitude in the partition function estimate for the initial reward
Here we compute the variance of the $\log{\hat{Z}(x)}$ estimate for the general case of using the initial reward for the policy fitting regression objective. Recall that the partition function can be expressed as a moment generating function for the distribution of the reward used in the objective 1 (see Appendix D for the derivation):
$Z(x)=\mathbb{E}_{r\sim P_{\mathrm{ref}}(\cdot\mid x)}\left[\exp\Bigl(\tfrac{1}{\beta}r\Bigr)\right],$
@@ Appendix I Noise amplitude in the partition function estimate for the initial reward
Appendix J Noise amplitude in the transformed reward estimate
@@
Appendix K Equivalence of all quantile transformation functions strictly increasing and upper-bounded on the interval $[0,1]$
In this section, we show that when $\beta$ is sufficiently small and the initial reward $\mathcal{R}$ can be approximated as a continuous variable, the optimal policy for a transformed reward $\mathbf{\tilde{\mathcal{R}}}(x,y)=f\left(\mathbf{\mathcal{R}}_{q}(x,y)\right)$ remains invariant under any choice of a strictly increasing continuous transformation function $f(t)$ defined on $[0,1]$ with finite values and a non-zero left derivative at t=1t=1.444Functions with zero left derivative at $t=1$ correspond to another equivalence class. More formally, it can be shown that there exist classes of equivalence. The $n$-th class consists of functions whose first $n$ left derivatives are all zero at $t=1$ but have a non-zero $n+1$ left derivative at $t=1$. We make a derivation only for the case of $n=0$, i.e., non-zero first left derivative. Others can be done similarly. Formally, as $\beta$ decreases, the policy converges to the same solution regardless of which function $f$ in this family is used.
@@
$\tilde{\mathcal{R}}(x,y)=f\left(\mathbf{\mathcal{R}}_{q}(x,y)\right),$
@@
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\times\left\{\frac{1}{\beta}f\left(F_{ref}(x,r)\right)-\frac{1}{\beta}g\left(F_{ref}(x,r)\right)+\log{\frac{Z_{g}(x)}{Z_{f}(x)}}\right\}$
@@
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\times\underbrace{\sum_{y:\mathcal{R}(x,y)=r}\pi_{ref}(y\mid x)}_{P_{ref}(r|x)}$
@@
where we used the fact that $f(1)=g(1)=0$ and $f^{\prime}_{-}(1)=g^{\prime}_{-}(1)=1$ after rescale and shift described in Appendix K.1.
@@
The partition function has the following asymptotics (see Appendix L for more details):
@@
Appendix L Asymptotics of the partition function $Z$ for the family of quantile transformation functions strictly increasing and upper-bounded on the interval $[0,1]$
@@
We assume mild regularity conditions, specifically $f\in C^{\infty}[0,1]$. We also assume that $f$ is normalized according to the rescale and shift procedure described in Appendix K.1.
Recall that the partition function can be expressed as a moment generating function for the distribution of the reward used in the objective 1 (see Appendix D for the derivation). Here we consider objective 1 with the generalized transformed reward $\tilde{\mathcal{R}}$:
[…]
where $\tilde{P}_{\mathrm{ref}}(r\mid x)$ denotes the probability distribution of the generalized transformed reward $\tilde{\mathcal{R}}(x,y)$ induced by the reference policy $y\sim\pi_{ref}(\cdot\mid x)$.
@@
$\tilde{Z}(x)=\mathbb{E}_{\tilde{r}\sim\tilde{P}_{\mathrm{ref}}(\cdot\mid x)}$ $\left[\exp\Bigl(\tfrac{1}{\beta}\tilde{r}\Bigr)\right]=\mathbb{E}_{r_{q}\sim\mathrm{U}[0,1]}\left[\exp\Bigl(\tfrac{1}{\beta}f\bigl(r_{q}\bigl)\Bigr)\right]=\int_{0}^{1}\exp{\left(\frac{1}{\beta}f\left(z\right)\right)}~dz.$
Here we used the fact that the probability distribution of the quantile-reward $\mathcal{R}_{q}(x,y)$ induced by the reference policy $y\sim\pi_{ref}(\cdot\mid x)$ is uniform from 0 to 1 (see Appendix E for the derivation).
@@
Note, that we assume the function $f$ to be normalized according to the rescale and shift procedure described in Appendix K.1 (as we require this property in the transformation functions equivalence proof K). Without this normalization the asymptotics will have a different form which can be derived similarly using the Laplace’s method.
Appendix M Detailed experimental protocol
@@
Llama 8B Tülu 3 SFT (lambert2024t) Llama-3.1-8B Tülu 3 SFT dataset
Mistral 7B Instruct v0.2 (jiang2024identifying) Mistral-7B-v0.2 Unknown public instruction datasets
@@
Magpie-Air (xu_magpie_2024) General Chat 98,000 2,000 Length >> 2048 tokens << 1%
UltraFeedback (cui_ultrafeedback_2024) General Chat 61,135 2,000 Length >> 2048 tokens << 1%
LeetCode (xia2025leetcodedataset) Coding 2,641 228 Length >> 2048 tokens or << 20 test cases << 3%
@@
General Chat ArmoRM (wang_interpretable_2024) Reward model with sequences truncated at 2048 tokens
General Chat Length-controlled ArmoRM (new) Length bias canceled via regression as in AlpacaEval 2 (dubois2024length)
@@
We release a reference implementation of QRPO based on a HuggingFace TRL  (vonwerra2022trl) trainer also supporting the baselines we trained (DPO, REBEL, SimPO) at https://github.com/CLAIRE-Labo/quantile-reward-policy-optimization. The codebase is based on the Python Machine Learning Research Template (Moalla_Python_Machine_Learning). It contains
@@
M.2 Initial SFT-only Models
We consider popular middle-scale models that have been instruction fine-tuned (IFT/SFT) but have not undergone RL fine-tuning. [^5] This is important as our experiments emulate the RL fine-tuning stage after the supervised fine-tuning stage in a post-training pipeline, and more importantly because reporting improvements on only specific benchmarks over already (usually closed-source) RL fine-tuned models is a biased view on the performance profile of these models which have been carefully tuned to balance multiple benchmarks. E.g., taking Llama-3-8B-Instruct and improving its AlpacaEval 2 scores by RL fine-tuning it on a narrow chat dataset is not a faithful way to report an algorithmic improvement as it does not show where the model has regressed. We choose allenai/Llama-3.1-Tulu-3-8B-SFT (Llama 8B Tülu 3 SFT) (lambert2024t), a fine-tuning of the meta-llama/Llama-3.1-8B base model on the Tülu 3 SFT dataset (lambert2024t), and mistralai/Mistral-7B-Instruct-v0.2 (Mistral 7B Instruct v0.2) (jiang2024identifying).
M.3 Datasets
For the general chat task, we use two popular alignment datasets, extensively used in previous work (meng2024simpo; gao_rebel_2024) consisting primarily of synthetic data.
Magpie-Align/Magpie-Air-DPO-100K-v0.1 (Magpie-Air) (xu_magpie_2024) is a strong alignment dataset containing 98000 training samples (and 2000 testing samples) with synthetic instructions and completions generated by Llama-3-8B-Instruct targeting mainly information seeking, and other categories such as creative writing, advice seeking, planning, and math. The chosen (resp. rejected) completions have been selected by picking the best (resp. worse) completions in a set of 5 completions rated by the RLHFlow/ArmoRM-Llama3-8B-v0.1 reward.
HuggingFaceH4/ultrafeedback_binarized (UltraFeedback) (cui_ultrafeedback_2024) consists of 61135 training samples (and 2000 testing samples) with instructions targeting instruction following, truthfulness, honesty, helpfulness, and completions generated by a large pool of models with different capabilities (between 7B and 70B open models and closed models including GPT4), annotated by GPT-4. This is the pre-processed version of the original UltraFeedback dataset that was used to train Zephyr-7B-$\beta$ (tunstall2023zephyr), where for each prompt, the completion with the highest overall score is taken as the chosen completion, and one of the remaining three at random as the rejected one.
For the coding task, we use newfacade/LeetCodeDataset (LeetCode) (xia2025leetcodedataset), which contains 2641 training samples (and 228 testing samples) comprising easy, medium, and hard LeetCode problems. Each problem has a median of 100 test cases and contains a correct solution that we use for additional SFT before performing the RL fine-tuning phase.
@@
M.4 Rewards
For the general chat task we use the RLHFlow/ArmoRM-Llama3-8B-v0.1 (ArmoRM) (wang_interpretable_2024) reward model with a maximum sequence length of $2048$. It is used to compute the rewards directly used by REBEL and QRPO in all distribution settings and to label the chosen and rejected preferences for DPO and SimPO in the off-policy setting. This also holds for Magpie-Air in the offline setting as its chosen and rejected annotations have been annotated with ArmoRM as well by xu_magpie_2024. For UltraFeedback the offline chosen and rejected annotations are not annotated by the ArmoRM, but most of the best models for the baselines turned out to be from the off-policy setting due to the low quality of the UltraFeedback offline completions.
Although ArmoRM has a verbosity component to reduce length bias, it still suffers from out-of-distribution length bias. i.e., its length bias has been minimized on specific outputs (original UltraFeedback), but is still present when computing the rewards on chats that diverge from the UltraFeedback outputs, which we did observe in our experiments. We therefore also report a length-controlled reward whose coefficients are computed for each model based on a regression that cancels the length component as done in AlpacaEval 2 (dubois2024length) (see Appendix M.5 for more details on the length-controlled reward computation).
For LeetCode, we use a Python sandbox for executing LLM-generated code in a safe environment with exclusive and controlled hardware limits, which we built for this project. The reward for a solution to a problem is the average number of test cases passed in the suite of tests for the problem. Each problem has a minimum of 20 tests and a median of 100 tests, and we cap the number of tests to run for each problem to the first 100 tests for computational efficiency. We use this average pass rate reward instead of a binary reward indicating whether all tests passed for two main reasons: (i) because the return value of the solution for edge cases is often not specified, still those appear in the test cases with arbitrary targets causing high rewards to collapse when binarized (e.g., passing the test cases with a rate of 0.99 because of a single failed test on an under-specified edge case would be converted to a 0 binary reward). (ii) because we assume a reward that can be treated as continuous in the theory of QRPO. [^6] We have also derived a version of QRPO for rewards with a small number of possible values in Appendix F which would be compatible with the binary reward in this case. We defer this setting to future work.
M.5 Length-controlled reward (LC-Reward) computation